BCAL Beam Tests 2012
- 1 Objectives
- 2 Coordination
- 3 Documentation
- 4 Logistics
- 5 Calculations
- 6 Simulations
- 7 Photos and Plots
- 8 E and T Counter energy boundaries
- 9 DAQ and Trigger
- 10 Run Plan
The possibility to carry out a BCAL beam test in the spring of 2012 has been discussed within the Calorimetry Working Group (CWG). Below, information is being gathered to agree upon the goals for this test before proceeding further and requesting parasitic time in Hall-B.
The central objective is to fully instrument a BCAL module with SiPMs, complete with their board electronics, cooling and mechanical assembly, coupled to Flash ADCs and readout by the planned Hall-D DAQ system. This would afford a comprehensive tryout of all aspects connected with the BCAL readout in a realistic beam environment and expose any issues that would require corrections or adjustments, before instrumenting the BCAL modules in the solenoid.
In no particular order, below are possible goals, for discussion.
- Instrument a module with 80 SiPMs, 40 per side. If possible, these should be instrumented as planned for the physics runs, with
- final lightguides
- pre-amplifier boards
- full cooling
- mechanical assemblies
- Flash ADCs and F1 TDCs
- the Hall-D DAQ system
- Data runs will be converted for analysis by ROOT.
- Such a test will allow verification of operation in a beam environment, and would give us valuable information in validating our MC studies on SiPM noise, thresholds, etc.
- If possible, a broad energy range should be used, from at least 500 MeV to 2 GeV. This would extend the range acquired during the 2006 beam tests, and would allow us to get a better handle on the floor term in the energy resolution, by going to higher energies.
- The module should be oriented to allow both 90 degree and shallow angle (>20 degrees) measurements. These would then be compared to the 2006 data and show the difference in behaviour between SiPMs and PMTs.
- To achieve a large energy range, the module may have to be placed at a location behind the tagger, where a full-length module would not fit. If this is the case, a shorter module can be used out of an existing prototype piece, or especially constructed in Regina following the completion of the 49 modules (48 plus the spare). Calculations would have to be carried out to determine the appropriate length so as not to clip the showers for the shallow angle measurements.
- If available by that time, Athens LEDs should be installed at least on two opposing BCAL cells, located centrally and surrounded by other cells so as to investigate operation as well as optical cross talk. However, as this stands to introduce additional complexity into an already complex setup, it is perhaps best left to be done on the bench, possible during cosmic tests.
The coordinators for this effort are Zisis and Elton. Information from mechanical (Tim) and electronics/readout (Fernando) will be coordinated through Elton. Eugene, Elton and Sascha will interact with Hall B folks to ensure a smooth procedure from hall access request to installation and running.
At the appropriate times, a request will be submitted to the Hall B Manager, followed by all other necessary documentation for running and safety. See the 2006 runs for reference.
- Elton's presentation to Hall B, Monday Feb 13.
- Proposal to Hall B
- Work Control Document
- Bcal Elec/Cool assembly test with LED
Following a conference call between IU and UR on October 17, the IU summary follows, with additional information added from feedback by Matt and Elton on October 19.
- Test station: Optimal location for our study appears to be the Tagger Alcove, under the upstream end of the tagger magnet. The available volume between the floor and the Tagger ranges from 2-3m down to 0m where the Tagger contacts the floor. The allowable width for any setup is about 1m wide. For this reasons and the fact that there is no crane access there, a small baby-cal, around 50cm will work, allowing for space to mount the readout. Elton notes that In the alcove, you only get about 5-15% of the incident beam energy. For higher energies, one needs to locate the detector down onto the Hall B floor where we may reach 30% of the energy. The space there is somewhat more crowded -- although it is being cleaned -- but should allow room for the BabyCal.
- Energy: The FCAL beam test will be planned to allow a scanning of a 20x20cm2 area of the device. IU expects the electron energy range of 5-40% of beam energy to strike the FCAL. This would also be handy for the BCAL, namely a scan would allow better extraction of the energy resolution floor term.
- Running mode: IU plans to use their own tagging scintillators to run entirely independent of the tagger, i.e. parasitically. UR will do the same.
- Gain Balancing: The IU group plans to centre their own tagged beam in each crystal for gain balancing. UR can do this using cosmics and electrons.
- Count Rate: The Tagger MOR runs at 20 MHz. IU expects 1-10 kHz in each of their 4"x4" blocks. This is manageable for the BCAL as well.
- Orientation: It may be advantageous to orient the BabyCal perpendicular to the Hall B beam line so as to intercept only a slice of the tagger rate, that over the width (~8-12cm) of the BabyCal. The beam envelope is expected to be ~2cm. We need to ensure that we do not clip the showers for small angles of incidence. We may consider adding veto counters to cover certain parts of the BabyCal. This requires further thinking and discussion.
- Positioner: this has been constructed and shipped to JLab. Its manual is available here.
- Electronics: Sascha Somov has been contacted to see what is involved in getting the electronic, ADCs, TDCs, discsiminators, crates, trigger and DAQ/online set up.
- from John, February 1: 2 pass photon until Feb 13; down few days to switch to electron running 1-2 weeks data with electrons; few days down to go back to photons; early to mid March few days down to install target then 3 pass for few weeks, then 5 pass. After mid-march there is no scheduled down period other than weekly beam studies on Wednesdays.
- The revised accelerator operations schedule is firm through May 2012. March and April look to be the most suitable for the beam test.
- Deadlines: The Flash ADC RFP will go out in early November whereas the electronics boards for the SiPMs need to be ordered by mid December. Mechanical drawings for the Readout Assembly are being made and should be done by mid December. First Article light guides (80 units) will be available before the end of the year. Therefore, all equipment should be available in principle for a beam test taking place possibly as soon as February, but more probably in March or April.
- Information from Kei:
- 2011 Hall B FCAL Beam Test
- Presentation at Collaboration Meeting
- Beamtest Photos and Plots
- EltonPictures.zip: pictures that Elton and Sasha took.
- HallBtaggerGeom.xls: Excel file created by Richard with dimensions of the hall.
- 66840-05050-05.pdf: Engineering drawing of the hall is
- ROOTplots.pdf: Plots created with a ROOT script.
- tagger/ Pictures of the tagger area.
- PS/ Pictures of the Pair Spectrometer area.
- beamdump/ Pictures of the beam dump area.
- proposal/ The proposal document is also on the docDB.
Initial simulations have been done by Irina and Andrei. We will resume these.
Photos and Plots
- Mini Cal positioner
- Light through full complement of light guides on one side, iPhone light shinning through the middle guides. Media:C1010003.JPG
- Second side of light guides glued onto mini-Bcal Media:C1010008.JPG
- Scot standing next to completed module except for one "K" guide, which has not yet arrived Media:C1010013.JPG
- Bcal electronics/cooling package: Front view with 40 SiPMs Media:A1010003.JPG
- Bcal electronics/cooling package: Side view Media:A1010004.JPG
- Bcal electronics/cooling package: Back side with output connectors Media:A1010005.JPG
- Installing electronics/cooling assembly onto mini-Bcal Media:Baby_BCal_Assembly_084.jpg
- mini-Bcal with electronics/cooling assemblies installed 1 Media:Baby_BCal_Assembly_087.jpg
- mini-Bcal with electronics/cooling assemblies installed 2 Media:Baby_BCal_Assembly_088.jpg
- Hall B - mini-Bcal at normal incidence Media:HallB_miniBcal_normal.JPG
- Hall B - mini-Bcal 22% point at 20 degrees to beam (front side) Media:HallB_miniBcal_20degFront.JPG
- Hall B - mini-Bcal 22% point at 20 degrees to beam (back side) Media:HallB_miniBcal_20degBack.JPG
- Hall B - mini-Bcal 35% point at 5 degrees to beam (view 1) Media:HallB_miniBcal_5degview1.JPG
- Hall B - mini-Bcal 35% point at 5 degrees to beam (view 2) Media:HallB_miniBcal_5degview2.JPG
- Hall B - mini-Bcal chiller Media:HallB_chiller.JPG
- Hall B - mini-Bcal Area under tagger, Fcal 22% point, Bcal 35% point Media:HallB_tagger_area.JPG
E and T Counter energy boundaries
DAQ and Trigger
- Trigger and Readout Electronics
- How to run BCAL DAQ
- Media:HallB beamtest periods.pdf
- One page summary of Bcal running periods.
- Beam current in nA as a function of time for the BCAL beam test in Hall B 2012
- Energies, angles of detector and various hand anotations
- Event Analyzer
- Rate Monitor
The initial checking of the Mini Cal test module will be attempted on the bench and possibly using cosmics in the EEL lab.
Check the Hall D E-log for updates.
- Following bench tests of boards by Fernando and SiPMs by Carl/Yi, examine operation of entire setup. Power up SiPMs and electronics. Look at signals on scope and then DAQ.
- Power up LED boards. Look at signal on scope and DAQ. Check trigger frequency and pulse width, as well as far/near ratio.
- Using cosmics, the SiPMs should be examined. In a few days in EEL 126, statistics may be an issue, but attempt to accumulate some data nonetheless.
SiPM Gain Inspection
- Based on the selected (anti-binned) SiPMs, examine and record their pulse heights on a scope. Discuss how to handle the ones summed in towers.
SiPM Bias Settings
4/10/12 Updated Bias Voltage Settings taking into account automatic bias adjustment by the thermistors on board
The first order correction due to the non-linear response of the thermistors is dV/dT = 0.006 V/deg C
- B100=74.280 V (T=25), 74.160 (T=5)
- B101=74.796 V (T=25), 74.676 (T=5)
- B102=75.308 V (T=25), 75.188 (T=5)
- B103=75.808 V (T=25), 75.688 (T=5)
- B104=74.791 V (T=25), 74.671 (T=5)
- B105=75.319 V (T=25), 74.199 (T=5)
- B106=75.290 V (T=25), 75.170 (T=5)
- B107=75.802 V (T=25), 75.682 (T=5)
3/27/12 ES. Based on the test voltage measurements on the board, the following bias voltages are suggested (assuming DeltaV/DeltaT = 0.056 V/deg C)
- B100=74.280 V (T=25), 74.045 (T=20.8), 73.877 (T-17.8)
- B101=74.796 V (T=25), 74.561 (T=20.8), 74.393 (T=17.8)
- B102=75.308 V (T=25), 75.073 (T=20.8), 74.905 (T=17.8)
- B103=75.808 V (T=25), 75.573 (T=20.8), 75.405 (T=17.8)
- B104=74.791 V (T=25), 74.556 (T=20.8), 74.388 (T=17.8)
- B105=75.319 V (T=25), 74.084 (T=20.8). 74.916 (T=17.8)
- B106=75.290 V (T=25), 75.055 (T=20.8), 74.887 (T=17.8)
- B107=75.802 V (T=25), 75.567 (T=20.8), 75.399 (T=17.8)
Cosmic Flux and Count Rate
The area of each cosmics counter is XX x YY = ZZZcm2. The cosmic flux crossing a unit horizontal area is 180 (130 hard/muons, 50 soft/electrons) m-2 s-1. This then leads to an expected rate for the hard muons of MMM s-1, or MMMM hr-1, or when divided equally among 4 columns, we get NNN hr-1 PMT-1. Thus, each cosmic run needs to be about QQQ hrs long to collect adequate statistics for an accurate determination of the ADC centroid and pedestals.
Main Experiment Goals
- Begin at 22.8% E/E0 point for 5-pass beam
- Re-verify operation of all SiPMs and LEDs once installed under the tagger.
- Re-check electronics, DAQ and online acquisition.
- Check SIPM gain balance. Adjust bias if necessary during restricted access.
- Check that Flashes and TDCs are seeing data through the online software.
- Activate LED monitoring system at recommended 12 V bias. Tune and see if appropriate, otherwise adjust.
- Take data. Start pipe lining to Regina for analysis once nominal data taking is established.
- Begin at the center of the module at normal incidence. Stay for a week or long enough to make sure data is of good quality.
- The first Wednesday after that, during maintenance, access Hall B and rotate module to 20 degrees.
- Take data at 20 degree incidence, 3cm from edge of Mini Cal. Showers will develop diagonally through 44X0. Stay at this point for a week or long enough to make sure data is of good quality.
- If 3-pass beam is delivered, repeat 90 and 20 degree measurements at the 22.8% E/E0 point.
- If time permits, repeat previous steps at 5% E/E0 point with 5- or 3-pass beam; whichever is available.
- Through all this period, change LED monitoring system bias to lower and higher values to get different calibration points.
- Establish appropriate ROOT trees from Analyzer.
- Investigate SiPM Flash ADC signals and TDC timing for consistency.
- Extract LED monitoring of near/far ends and cell-to-cell for all used bias values.
- Extract the energy and timing resolution and number of photo electrons for all data runs in the analysis.